# Zero Trust Architecture: A Systematic Literature Review

### v1

@@ 5.1.6 Virtual Environments
Cloud computing is increasingly being adopted by companies and organizations as it provides flexibility, scalability, and cost efficiency [126]. However, with the amounts of data being managed in and stored on cloud platforms growing at breakneck pace, protecting sensitive data and its confidentiality against unauthorized access and use is critical [127].The Okta breach, which involved an identity and access management company, is an example of a recent cloud security incident. The attackers used compromised credentials to sneak into Okta’s systems and obtain cookies and sensitive session tokens via HTTP Archive or HAR files [128]. The increased level of complexity in cloud platforms makes it challenging for system administrators as well as the conventional security mechanisms to identify and mitigate security threats from adversaries [74]. To effectively secure the cloud infrastructure, a ZT-based model has been proposed by reasearchers which provides a dynamic approach to securing the cloud infrastructure which might not be possible through PBSM [5].
@@
Several articles have highlighted the feasibility and potential of ZTA in securing cloud and edge platforms. In [74], Changhee et al. proposed a framework called S-ZAC (”Service Mesh Zero Trust Acces Control”) to enhance the security of service mesh control plane through SGX technology in order to ensure the CIA of information even in the presence of privilege attackers like malicious cloud service provider. By implementing a prototype of the S-ZAC, the authors evaluate its performance and effectiveness in enforcing access control and provisioning within cloud environments, ultimately aiming to establish a secure ZTA for cloud-based applications. Another ZTA framework for cloud security is presented in [127] where the authors show a careful integration of encryption, Blockchain technology, sharding concept, and Zero Knowledge Proofs (ZKPs) to create a secure cloud environment for data storage and migration. The proposed framework adopts a ZT approach, ensuring that server handlers do not have direct access to the raw data, and thus enhancing overall data security and privacy throughout the cloud environment. To secure the edge platform which are used due to the limited resources of IoT devices, Cem et al. [129] proposed a blockchain-based ZTA framework. The proposed framework consists of three main component groups: ZTA components which enforce policies and manage access requests, blockchain components which serves as an immutable database, storing request history and facilitating trustworthiness verification through a consensus mechanism. , and IoT components which comprise stationary IoT devices with sensors and users. Through experimental processes, the authors evaluate the framework in terms of performance, complexity, and scalability, and indicate its applicability in smart cities contexts.
@@
On another perspective, Jyoti and Rajeshkumar [126] and Muhammad et al. [5] conducted surveys to investigate the usage of cloud in implementing ZT security strategies. In [126], the authors explore the relevance of establishing trust in cloud and how it compares between traditional approach and ZT approach, security challenges in cloud, and recommendation for implementing ZT in cloud. Similarly, in [5], the authors considered cloud as a platform to implement ZTA and reviewed several existing works that explore it in their research.
@@
The metaverse represents a convergence of virtual and augmented reality, creating shared, immersive spaces for gaming, social interaction, and commerce [130]. Its reliance on real-time data exchange, digital identities, and asset transactions introduces significant security concerns, including identity theft, unauthorized access, and data breaches [131], [132]. However, traditional security frameworks that rely on perimeter-based defences are not performing effectively in these open and dynamic digital environments [130]. To address this security concern of the virtual environment such as metaverse, in [133], the research explores the feasibility of implementing a blockchain-based ZTA framework that enable secure and transparent interaction between user and applications in Metaverse environment. The reviews various security issues and presented three case studies that hint on the practicality of blockchain-ZTA solution for the Metaverse environment. Similarly, in [134], while indicating the issue of impersonation and user authentication challenges in virtual reality (VR) such as the Metaverse, Ruizhi et al. conducted a research that investigated four factors regarding the security of VR which include effectiveness of using federated learning (FL) for ensuring privacy of user biometric data, and improving the usability of ZT security with adaptive VR-based authentication. The study concluded that traditional FL algorithms are not quite suitable for authentication of VR users using biometrics which leads to a less accuracy. They indicated that each user’s local data represents only their legitimate or authenticated interactions (positive label data), without including any examples of failed, incorrect, or unauthorized interactions (negative label data).
Similarly, Ikram [62] proposed a blockchain-enabled ZTA framework which integrates decentralized identity verification through blockchain and other techniques such as cryptography, graph theory, and machine learning. The authors demonstrate the effectiveness of their proposed framework and evaluate it based on six metrics among which are dynamic response to threats, decentralized authentication and response, and scalability. Also, Asma et al. [135] integrated device password checks, device authentication, and hand behavior analysis in their proposed ZT framework for securing the Metaverse. This setup grant initial access only to authorized users, after which user’s voice is continuously analyzed in real-time to confirm their identity and determine their access rights based on three trust levels.
@@ 5.1.7 Transport and Logistics
The transport and logistics sector plays a crucial role in global trade and commerce, relying heavily on interconnected systems and technologies. With the increasing adoption of digital transformation and IoT in this domain, implementing Zero Trust Architecture (ZTA) is critical for ensuring robust security, ensures compliance with regulatory standards, and increases operational resilience against evolving cyber threats. Key applications areas identified in the reviewed articles where ZTA is implemented include supply chain, vehicular communication systems, and FPGA supply chain.
@@
The supply chain involves complex interactions among manufacturers, suppliers, distributors, and retailers, often across international boundaries. While these entities can be dispersed around, the interaction can also be complex which makes the system to be vulnerable and exposed to security issues that need to be properly addressed. In 2023, a ransomware attack on one of the suppliers of Applied Materials, semi-conductor giant company, disrupted the supply chain, causing shipment delays and financial losses estimated at up to \$250 million [136]. To address security vulnerabilities like this, Zachary and Joseph [137] proposed a ZT supply chain security framework. The authors highlighted the porous and poorly defined perimeter of the supply chain industry which enables malicious entities to intercept sensitive information and interfere with operations. They examine how an enterprise could move towards ZT, map ZT principles to the supply chain, and outline a research plan by looking at ZT from the perspectives of various organisational theories. Similarly, a report in [138] indicates that the DoD has adopted ZT approach in buying microelectronics through an executive order released in 2021.
@@
Unmanned Aerial Vehicles (UAV) or drones are widely used in logistics for delivering goods, monitoring supply chain facilities, and surveillance. These devices often communicates with, and are being controlled by a control station. Possible attack scenarios include intercepting data between UAVs and control stations, spoofing attacks, and integrating unauthorized drones into the logistics network. To assist in mitigating security threats, Ekramul Haque et al. [139] proposed a deep learning driven ZT framework for improving UAV security. Through continuous authentication strategy which is conducted at a regular time interval of 10 seconds, the framework distinguishes between UAV types based on their properties and behaviour using radio frequency (RF) signals. The proposed models demonstrate an appreciable performance in enhancing the security of the UAV systems. Similarly, in [69], the authors proposed a decentralized ZTA framework that ensure the security and authenticity of transmitted data and UAV swarms. They used federated learning to maintain continuous trust evaluation and increased levels of access controls, and Blockchain technology to ensure the trustworthiness of UAV nodes based on security, privacy, and reliability. According to experimental findings, the suggested architecture can provide UAVs with high levels of authenticity and intrusion prevention.
@@
Furthermore, a lightweight ZTA framework is proposed in [80] to enhance security for vehicle-to-infrastructure (V2I) and vehicle-to-vehicle (V2V) communication in 5G vehicular ad hoc networks (VANETs). The framework ensure continuous authentication, monitoring, and fine-grained access control through the integration of four key components - a central server that manages identity verification, access control, monitoring, and authenticate all vehicles through MFA using an OTP and digital signature before they can send messages, micro-segmenting the VANETs into secure clusters to limit lateral movement of threat actors, anomaly detection capability through many security checks, and encrypting the communication between V2V and V2I. Experimental results show that the framework has high throughput and can protect against impersonation attack while ensuring secure vehicular communication. Similarly, while leveraging Blockchain technology, Chengzu et al. proposed two ZTA frameworks in separate articles to secure UAV delivery system [140], [141]. They highlighted the risks associated with UAV delivery process where the UAVs need to store or have access to sensitive information about the transmitted parcel and the receiver which can be exposed to spoofing or injection attacks. In [140], MFA and Self-Sovereign Identity (SSI) techniques are used to ensure continuous authentication and privacy of receiver’s data, while in [141], homomorphic encryption is used. Both frameworks are shown to demonstrate good performance in addressing the identified security vulnerabilities through experimental evaluation.
@@
In addition, to protect Internet of Vehicle Things (IoVT) against intrusions which can result in data leakage and hijacking of vehicle control by attackers, a ZTA framework is proposed in [75] which combines attribute-based access control (ABAC) and trust evaluation. The framework utilizes attributes for access control and incorporates user trust scores to enhance security and reliability. The introduction of trust score calculation and evaluation as an integral part of the access control process addresses the limitation of ABAC in fully capturing user intentions and enhances the security and flexibility of authorization. Experimental results show the superior performance and effectiveness of the framework compared with conventional authentication and access control schemes.
@@
Field-Programmable Gate Arrays (FPGAs) are essential components in transport and logistics systems, often used for real-time processing in automated vehicles and IoT devices. Participants from around the globe work together in the contemporary FPGA supply chain to guarantee its smooth operation [142]. Nonetheless, this leads to a number of challenges and possible problems, such as a malicious third-party vendor tampering with the actual intellectual property (IP) core [143], misconduct by design house employees like changing the IC design or stealing and leaking IP [144], the introduction of hardware Trojans during the manufacturing process, and the unapproved overproduction of ICs [145]. To mitigate some of these issues, Akshay et al. [61] proposed a ZTA framework using key ZT principles like access control, authorisation, and continuous monitoring, as well as enabling technologies like blockchain for user verification and multi-factor authentication and ring oscillator physical unclonable functions (ROPUFs) for device authentication. The framework eliminates blind trust and guarantees that every element using or requesting access to the resources of an FPGA supply chain goes through a rigorous verification process. To assess the framework’s efficacy, every ZT principle is investigated using a different attack scenario.
@@ 5.1.8 Military and Defense
Defense and tactical systems require stringent security measures to protect critical operations and sensitive data in dynamic environments. ZTA addresses these needs through continuous verification, micro-segmentation, and secure inter-domain collaboration, thus, ensuring resilience against evolving cybersecurity threats. From the reviewed articles, Alexandre et al. [147] proposes a ZTA framework that will allow different nations and inter-domains to collaborate in a tactical mission and share resources without trusting each other. The framework ensures secure and authorized access to data and services where every access request is explicitly verified without the need a for federation partner to trust the other. While each domain operates independently and controls its resources and infrastructure, the framework integrates a component called ”Remote Attestation Verifier” which is deployed at the resource-requesting domain but managed by the resource domain. This approach ensures secure interoperability between domains without relying on implicit trust, enabling the seamless and secure sharing of resources while upholding the stringent security requirements of ZT environments.
Similarly, Youngho et al. [25] explores ZTA as a dynamic cybersecurity framework for defense, emphasizing continuous verification and security contexts rather than assuming trust based on network location, core challenge areas in defense cybersecurity and proposes technical solutions such as Multi-Factor Authentication (MFA) for mitigating identity fabrication and credential misuse; logical network segmentation using Software-Defined Networking (SDN) and micro-segmentation to address network security and monitoring issues; and data-driven approaches, including User and Entity Behavior Analytics (UEBA), for detecting insider threats and malicious activities. Together, these methodologies enable greater resilience against today’s online threats and meet key security requirements.
@@
Furthermore, Alanoud and Abdullah [66], though an SLR offer a framework that incorporates existing security best practices and emerging technologies to address the unique security requirements of military UAVs. They argue that as UAVs become increasingly integral to military operations, traditional PBSMs are no longer sufficient to protect them from sophisticated cyberattacks. Similarly, they indicated that securing UAVs involves addressing vulnerabilities across four distinct levels which are: protecting the physical components of the UAV from tampering or unauthorized access (Hardware Security), ensuring the integrity and security of the software and firmware that control the UAV’s operation (Software Security), Securing the communication channels between the UAV and the ground control station (GCS), as well as between UAVs in collaborative missions (Communication/Network Security), and Protecting the integrity of data collected by the UAV’s sensors to prevent manipulation or spoofing (Sensor Security). The study gives a detail analysis on the advantages, threats, and effectiveness of employing ZT solutions for UAV security.
@@ 5.1.9 Manufacturing
The manufacturing sector is increasingly leveraging digital transformation through technologies such as Industrial IoT (IIoT), smart factories, and automation. These advancements have improved efficiency and productivity but have also expanded the attack surface, exposing critical systems to sophisticated cyber threats. Some notable cyber-attacks that affected the manufacturing sector include the Norsk Hydro ransomware attack in 2019 [150] which targeted the Norwegian aluminum giant Norsk Hydro and forces the company to halt operations at several plants and revert to manual operations. The estimated financial loss exceeded \$71 million. Also, the Honda cyber-attack in 2020 [151] which disrupted Honda global operations, including manufacturing plants and customer services, and impacted several production lines leading to operational downtime and delays. These incidents demonstrated the risks associated with the connectivity of manufacturing plants and the need for robust security.
From the reviewed articles, ZTA is shown to offer solution by introducing stringent security measures tailored to the unique needs of the manufacturing sector. For example, in [152], Renlog et al. highlighted the increasing interaction and blurred security boundaries between ship platforms which are currently depending on static access control to secure sensitive resources. To address this challenge, they proposed a ZT framework with dynamic access control and trust evaluation for secure and fine-grained access permissions. In the framework, the identity of user and role are verified and illegal roles are intercepted and access denied while legitimate access are continuously evaluated. Similarly, emphasizing on the issue of single point vulnerability and the associated challenge of deploying existing frameworks due to lack of proper workflow design, a ZT framework for enhancing the security of smart factories is proposed in [153]. The distributed ZT framework was designed based on software-defined perimeter (SDP) and cloud-edge-gateway, and the simulation results show improved performance in terms of efficiency and security.

### v2

@@ 5.1.6 Virtual Environments
Cloud computing is increasingly being adopted by companies and organizations as it provides flexibility, scalability, and cost efficiency [132]. However, with the amounts of data being managed in and stored on cloud platforms growing at breakneck pace, protecting sensitive data and its confidentiality against unauthorized access and use is critical [133].The Okta breach, which involved an identity and access management company, is an example of a recent cloud security incident. The attackers used compromised credentials to sneak into Okta’s systems and obtain cookies and sensitive session tokens via HTTP Archive or HAR files [134]. The increased level of complexity in cloud platforms makes it challenging for system administrators as well as the conventional security mechanisms to identify and mitigate security threats from adversaries [80]. To effectively secure cloud infrastructure, a ZT-based model has been proposed by researchers which provides a dynamic approach to securing the cloud infrastructure which might not be possible through PBSM [5].
@@
Several articles have highlighted the feasibility and potential of ZTA in securing cloud and edge platforms. In [80], Changhee et al. proposed a framework called S-ZAC (”Service Mesh Zero Trust Acces Control”) to enhance the security of service mesh control plane through SGX technology in order to ensure the CIA of information even in the presence of privilege attackers like malicious cloud service provider. By implementing a prototype of the S-ZAC, the authors evaluate its performance and effectiveness in enforcing access control and provisioning within cloud environments, ultimately aiming to establish a secure ZTA for cloud-based applications. Another ZTA framework for cloud security is presented in [133] where the authors show a careful integration of encryption, Blockchain technology, sharding concept, and Zero Knowledge Proofs (ZKPs) to create a secure cloud environment for data storage and migration. The proposed framework adopts a ZT approach, ensuring that server handlers do not have direct access to the raw data, and thus enhancing overall data security and privacy throughout the cloud environment. To secure the edge platforms which are used due to the limited resources of IoT devices, Cem et al. [135] proposed a blockchain-based ZTA framework. The proposed framework consists of three main component groups: ZTA components which enforce policies and manage access requests, blockchain components which serves as an immutable database, storing request history and facilitating trustworthiness verification through a consensus mechanism. , and IoT components which comprise stationary IoT devices with sensors and users. Through experimental processes, the authors evaluated the framework in terms of performance, complexity, and scalability, and indicated its applicability in smart cities contexts.
@@
On another perspective, Jyoti and Rajeshkumar [132] and Muhammad et al. [5] conducted surveys to investigate the usage of cloud in implementing ZT security strategies. In [132], the authors explore the relevance of establishing trust in cloud and how it compares between the traditional approach and ZT approach, security challenges in cloud, and recommendation for implementing ZT in cloud. Similarly, in [5], the authors considered the cloud as a platform to implement ZTA and reviewed several existing works that explore it in their research.
@@
The metaverse represents a convergence of virtual and augmented reality, creating shared, immersive spaces for gaming, social interaction, and commerce [136]. Its reliance on real-time data exchange, digital identities, and asset transactions introduces significant security concerns, including identity theft, unauthorized access, and data breaches [137], [138]. However, traditional security frameworks that rely on perimeter-based defenses are not performing effectively in these open and dynamic digital environments [136]. To address this security concern of the virtual environment such as metaverse, in [139], the research explores the feasibility of implementing a blockchain-based ZTA framework that enables secure and transparent interaction between the user and applications in the Metaverse environment. The reviews various security issues and presented three case studies that hint on the practicality of blockchain-ZTA solution for the Metaverse environment. Similarly, in [140], while indicating the issue of impersonation and user authentication challenges in virtual reality (VR) such as the Metaverse, Ruizhi et al. conducted a research that investigated four factors regarding the security of VR which include effectiveness of using federated learning (FL) for ensuring privacy of user biometric data, and improving the usability of ZT security with adaptive VR-based authentication. The study concluded that traditional FL algorithms are not quite suitable for authentication of VR users using biometrics which leads to a less accuracy. They indicated that each user’s local data represents only their legitimate or authenticated interactions (positive label data), without including any examples of failed, incorrect, or unauthorized interactions (negative label data).
Similarly, Ikram [68] proposed a blockchain-enabled ZTA framework which integrates decentralized identity verification through blockchain and other techniques such as cryptography, graph theory, and machine learning. The authors demonstrate the effectiveness of their proposed framework and evaluate it based on six metrics, among which are dynamic response to threats, decentralized authentication and response, and scalability. Also, Asma et al. [141] integrated device password checks, device authentication, and hand behavior analysis in their proposed ZT framework for securing the Metaverse. This setup grant initial access only to authorized users, after which user’s voice is continuously analyzed in real-time to confirm their identity and determine their access rights based on three trust levels.
@@ 5.1.7 Transport and Logistics
The transport and logistics sector plays a crucial role in global trade and commerce, relying heavily on interconnected systems and technologies. With the increasing adoption of digital transformation and IoT in this domain, implementing Zero Trust Architecture (ZTA) is critical for ensuring robust security, compliance with regulatory standards, and increases operational resilience against evolving cyber threats. Key applications areas identified in the reviewed articles where ZTA is implemented include the supply chain, vehicular communication systems, and the FPGA supply chain.
@@
The supply chain involves complex interactions among manufacturers, suppliers, distributors, and retailers, often across international boundaries. While these entities can be dispersed around, the interaction can also be complex which makes the system to be vulnerable and exposed to security issues that need to be properly addressed. In 2023, a ransomware attack on one of the suppliers of Applied Materials, semi-conductor giant company, disrupted the supply chain, causing shipment delays and financial losses estimated at up to \$250 million [142]. To address security vulnerabilities like this, Zachary and Joseph [143] proposed a ZT supply chain security framework. The authors highlighted the porous and poorly defined perimeter of the supply chain industry which enables malicious entities to intercept sensitive information and interfere with operations. They examine how an enterprise could move towards ZT, map ZT principles to the supply chain, and outline a research plan by looking at ZT from the perspectives of various organizational theories. Similarly, a report in [144] indicates that the DoD has adopted the ZT approach in buying microelectronics through an executive order released in 2021.
@@
Unmanned Aerial Vehicles (UAV) or drones are widely used in logistics for delivering goods, monitoring supply chain facilities, and surveillance. These devices often communicate with, and are being controlled by a control station. Possible attack scenarios include intercepting data between UAVs and control stations, spoofing attacks, and integrating unauthorized drones into the logistics network. To assist in mitigating security threats, Ekramul Haque et al. [145] proposed a deep learning driven ZT framework for improving UAV security. Through continuous authentication strategy which is conducted at a regular time interval of 10 seconds, the framework distinguishes between UAV types based on their properties and behaviour using radio frequency (RF) signals. The proposed models demonstrate an appreciable performance in enhancing the security of the UAV systems. Similarly, in [75], the authors proposed a decentralized ZTA framework that ensures the security and authenticity of transmitted data and UAV swarms. They used federated learning to maintain continuous trust evaluation and increased levels of access controls, and Blockchain technology to ensure the trustworthiness of UAV nodes based on security, privacy, and reliability. According to experimental findings, the suggested architecture can provide UAVs with high levels of authenticity and intrusion prevention.
@@
Furthermore, a lightweight ZTA framework is proposed in [86] to enhance security for vehicle-to-infrastructure (V2I) and vehicle-to-vehicle (V2V) communication in 5G vehicular ad hoc networks (VANETs). The framework ensures continuous authentication, monitoring, and fine-grained access control through the integration of four key components - a central server that manages identity verification, access control, monitoring, and authenticate all vehicles through MFA using an OTP and digital signature before they can send messages, micro-segmenting the VANETs into secure clusters to limit lateral movement of threat actors, anomaly detection capability through many security checks, and encrypting the communication between V2V and V2I. Experimental results show that the framework has high throughput and can protect against impersonation attacks while ensuring secure vehicular communication. Similarly, while leveraging Blockchain technology, Chengzu et al. proposed two ZTA frameworks in separate articles to secure UAV delivery system [146], [147]. They highlighted the risks associated with UAV delivery process where the UAVs need to store or have access to sensitive information about the transmitted parcel and the receiver which can be exposed to spoofing or injection attacks. In [146], MFA and Self-Sovereign Identity (SSI) techniques are used to ensure continuous authentication and privacy of receiver’s data, while in [147], homomorphic encryption is used. Both frameworks are shown to demonstrate good performance in addressing the identified security vulnerabilities through experimental evaluation.
@@
In addition, to protect the Internet of Vehicle Things (IoVT) against intrusions which can result in data leakage and hijacking of vehicle control by attackers, a ZTA framework is proposed in [81] which combines attribute-based access control (ABAC) and trust evaluation. The framework utilizes attributes for access control and incorporates user trust scores to enhance security and reliability. The introduction of trust score calculation and evaluation as an integral part of the access control process addresses the limitation of ABAC in fully capturing user intentions and enhances the security and flexibility of authorization. Experimental results show the superior performance and effectiveness of the framework compared with conventional authentication and access control schemes.
@@
Field-Programmable Gate Arrays (FPGAs) are essential components in transport and logistics systems, often used for real-time processing in automated vehicles and IoT devices. Participants from around the globe work together in the contemporary FPGA supply chain to guarantee its smooth operation [148]. Nonetheless, this leads to a number of challenges and possible problems, such as a malicious third-party vendor tampering with the actual intellectual property (IP) core [149], misconduct by design house employees like changing the IC design or stealing and leaking IP [150], the introduction of hardware Trojans during the manufacturing process, and the unapproved overproduction of ICs [151]. To mitigate some of these issues, Akshay et al. [67] proposed a ZTA framework using key ZT principles like access control, authorisation, and continuous monitoring, as well as enabling technologies like blockchain for user verification and multifactor authentication and ring oscillator physical unclonable functions (ROPUFs) for device authentication. The framework eliminates blind trust and guarantees that every element using or requesting access to the resources of an FPGA supply chain goes through a rigorous verification process. To assess the framework’s efficacy, every ZT principle is investigated using a different attack scenario.
@@ 5.1.8 Military and Defense
Defense and tactical systems require stringent security measures to protect critical operations and sensitive data in dynamic environments. ZTA addresses these needs through continuous verification, micro-segmentation, and secure inter-domain collaboration, thus, ensuring resilience against evolving cybersecurity threats. From the reviewed articles, Alexandre et al. [153] proposes a ZTA framework that will allow different nations and inter-domains to collaborate in a tactical mission and share resources without trusting each other. The framework ensures secure and authorized access to data and services where every access request is explicitly verified without the need a for federation partner to trust the other. While each domain operates independently and controls its resources and infrastructure, the framework integrates a component called ”Remote Attestation Verifier”, which is deployed in the resource-requesting domain, but is managed by the resource domain. This approach ensures secure interoperability between domains without relying on implicit trust, enabling the seamless and secure sharing of resources while upholding the stringent security requirements of ZT environments.
Similarly, Youngho et al. [25] explores ZTA as a dynamic cybersecurity framework for defense, emphasizing continuous verification and security contexts rather than assuming trust based on network location, core challenge areas in defense cybersecurity and proposes technical solutions such as multifactor Authentication (MFA) for mitigating identity fabrication and credential misuse; logical network segmentation using Software-Defined Networking (SDN) and micro-segmentation to address network security and monitoring issues; and data-driven approaches, including User and Entity Behavior Analytics (UEBA), for detecting insider threats and malicious activities. Together, these methodologies enable greater resilience against today’s online threats and meet key security requirements.
@@
Furthermore, Alanoud and Abdullah [72], through an SLR offer a framework that incorporates existing security best practices and emerging technologies to address the unique security requirements of military UAVs. They argue that as UAVs become increasingly integral to military operations, traditional PBSMs are no longer sufficient to protect them from sophisticated cyberattacks. Similarly, they indicated that securing UAVs involves addressing vulnerabilities across four distinct levels which are: protecting the physical components of the UAV from tampering or unauthorized access (Hardware Security), ensuring the integrity and security of the software and firmware that control the UAV’s operation (Software Security), Securing the communication channels between the UAV and the ground control station (GCS), as well as between UAVs in collaborative missions (Communication/Network Security), and Protecting the integrity of data collected by the UAV’s sensors to prevent manipulation or spoofing (Sensor Security). The study gives a detail analysis on the advantages, threats, and effectiveness of employing ZT solutions for UAV security.
@@ 5.1.9 Manufacturing
The manufacturing sector is increasingly leveraging digital transformation through technologies such as Industrial IoT (IIoT), smart factories, and automation. These advancements have improved efficiency and productivity but have also expanded the attack surface, exposing critical systems to sophisticated cyber threats. Some notable cyberattacks that affected the manufacturing sector include the Norsk Hydro ransomware attack in 2019 [156] which targeted the Norwegian aluminum giant Norsk Hydro and forces the company to halt operations at several plants and revert to manual operations. The estimated financial loss exceeded \$71 million. Also, the Honda cyberattack in 2020 [157] which disrupted Honda global operations, including manufacturing plants and customer services, and impacted several production lines leading to operational downtime and delays. These incidents demonstrated the risks associated with the connectivity of manufacturing plants and the need for robust security.
From the reviewed articles, ZTA is shown to offer solution by introducing stringent security measures tailored to the unique needs of the manufacturing sector. For example, in [158], Renlog et al. highlighted the increasing interaction and blurred security boundaries between ship platforms which are currently depending on static access control to secure sensitive resources. To address this challenge, they proposed a ZT framework with dynamic access control and trust evaluation for secure and fine-grained access permissions. In the framework, the identity of user and role are verified and illegal roles are intercepted and access denied while legitimate accesses are continuously evaluated. Similarly, emphasizing on the issue of single point vulnerability and the associated challenge of deploying existing frameworks due to lack of proper workflow design, a ZT framework for enhancing the security of smart factories is proposed in [159]. The distributed ZT framework was designed based on software-defined perimeter (SDP) and cloud-edge-gateway, and the simulation results show improved performance in terms of efficiency and security.